If it takes two men three hours to dig a hole 4 meter long, four meters wide, and four meters deep, how long would it take the same two men to dig a hole 12 meters long, 12 meters wide and 12 meters deep?
Answer by god2012(113)

total number of people = 2
Total duration = 3 hrs
Total volume of the hole = 4 * 4 * 4 = 64 cubic m
Required hole volume = 12 * 12 *12
Total time taken by 2 people = (12*12*12 / 64 ) * 3 = 81 hrs